Problem 2-6A Condensed balance sheet and income statement data for Sheffield Corp. are presented as follows Sheffield Corp Balance Sheets December 31 Assets Cash Receivables (net) Other current assets Long-term investments Property, plant, and equipment (net) Total assets Liabilities and Stockholder Current liabilities Long-term liabilities Common stock Retained earnings Total liabilities and stockholders equity 2017 2016 83,300 103,300 62,000 523,300 $803,040 $31,140 23,140 75,300 86,300 60,000 483,300 728,040 s Equity $78,140 73,140 103,300 313,300 238,300 $803,040 $728,040 93,300 343,300 288,300

Sheffield Corp. Income Statements For the Years Ended December 31 2017 2016 Sales revenue Cost of goods sold Operating expenses (including income taxes) Net income $778,000 $707,000 440,000400,000 240,000 220,000 $98,000 87,000 Additional information: Net cash from operating activities Cash used for capital expenditures Dividends paid on common shares Weighted-average number of shares outstanding $115,340 $59,140 $48,140 $38,000 $48,000 $18,140 30,000 33,000

Compute these values and ratios for 2016 and 2017. (Round Earnings per share to 2 decimal places, e.g. $2.78 and Current Ratio and Debt to assets ratio to 1 decimal place, e.g. 15.2. Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).) 2016 2017 (a) Earnings per share. (b) Working capital.$ (c) Current ratio. (d) Debt to assets ratio. (e) Free cash flow.

Anyone know these values and can explain how to get them?
Answer #1

Please provide rating... Let me know if you need any clarification

a) Earning per share
=Net income/weighted average number of share outstanding
2016 2017
Net income 87000 98000
Weighted average number of share outstanding 30000 33000
Earning per share          2.90                  2.97
b) Working cpital
=Current asset - Current liabilities
2016 2017
Cash 23140 31140
Receivable 75300 83300
Other current asset 86300 103300
Total current asset 184740 217740
Total current liabilities 73140 78140
Working capital = Current asset - Current liabilities 111600 139600
c) Current ratio
=Current asset / Current liabilities
2016 2017
Total current asset 184740 217740
Total current liabilities 73140 78140
Current ratio = Current asset / Current liabilities            2.5                     2.8
Answer = 2.5:1 2.8:1
d) Debt to asset ratio
=Total liabilities / Total asset
2016 2017
Total liabilities 176440 =73140+103300 171440 =78140+93300
Total asset 728040 803040
Debt to asset ratio = 24.2% 21.3%
e) Free cash flow
=Cash flow from operating activities - Capital expenditure - Dividend paid
2016 2017
Cash flow from operating activities 59140 115340
Less : Capital expenditure -38000 -48140
less : Dividend paid -18140 -48000
Free cash flow 3000 19200